OpenStudy (lostinmath101):

Sin20Cos80-Cos20Sin80

OpenStudy (campbell_st):

this is the expansion of sin(a - b) where a = 20 and b = 80 or sin(-60)

OpenStudy (lostinmath101):

I see..... thanks

OpenStudy (lostinmath101):

Still need some help

OpenStudy (anonymous):

check the question.....i think its wrong...it should be like this sin10cos80-cos10sin80 solving...... sin(90-80)cos 80-cos(90-80)sin 80=1/cos80*cos80-1/sin80*sin 80=1-1=0

OpenStudy (lostinmath101):

Well it says what I typed on the paper so......................

OpenStudy (inkyvoyd):

Sin20Cos80-Cos20Sin80 Sin (a-b)=sin a cos b-sin b cos a

OpenStudy (anonymous):

\[ \sin(-60^ \circ) = - \sin (60^\circ) = - \frac{\sqrt{3}} {2} \]